# S-Graphs 2.0 – A Hierarchical-Semantic Optimization and Loop Closure for SLAM

### v2

@@ Abstract
The hierarchical structure of 3D scene graphs shows a high relevance for representations purposes, as it fits common patterns from man-made environments. But, additionally, the semantic and geometric information in such hierarchical representations could be leveraged to speed up the optimization and management of map elements and robot poses, something that is not done by the methods in the literature.
@@
In this direction, we present our work Situational Graphs 2.0 (S-Graphs 2.0), which leverages the hierarchical structure of indoor scenes for efficient data management and optimization. Our algorithm begins by constructing a situational graph that represents the environment into four layers: Keyframes, Walls, Rooms, and Floors. Our first novelty lies in the front-end, which includes a floor detection module capable of identifying stairways and assigning floor-level semantic relations to the underlying layers. Floor-level semantics allows us to propose a floor-based loop closure strategy, that effectively rejects false positive closures that typically appear due to aliasing between different floors of a building. Our second novelty lies in leveraging our representation hierarchy in the optimization. Our proposal consists of: (1) local optimization over a window of recent keyframes and their connected components across the four representation layers, (2) floor-level global optimization, which focuses only on keyframes and their connections within the current floor during loop closures, and (3) room-level local optimization, marginalizing redundant keyframes that share observations within the room, which reduces the computational footprint. We validate our algorithm extensively in different real multi-floor environments. ‘ Our approach shows state-of-art-art accuracy metrics in large-scale multi-floor environments, estimating hierarchical representations up to faster, in average, than competing baselines. Our code is open-sourced at http://github.com/snt-arg/lidar_situational_graphs
Index Terms:
SLAM, Localization, MappingI Introduction
@@
SLAM, that has typically focused on geometric aspects, should also enable a robot with a high-level understanding of its environment [1]. The combination of 3D scene graphs with traditional SLAM graphs have recently sparked great interest given its compact and intuitive approach to representing an environment as a hierarchy of layers [2], [3]. In this direction, Situational Graphs (S-Graphs) [4] model the environment as a four-layered optimizable graph composed of of Keyframes, Walls, Rooms and Floors. As the robot explores large environments, however, the size of the graph grows, along with the its computational footprint. Essentially, while S-Graphs and 3D scene graphs are motivated by the need of higher-level representations, none of them have yet tailored their optimization to the specifics of these hierarchical representations. Leveraging them will result in a lower computational footprint, enabling real-time graph optimization for larger maps, which will result finally in enhanced accuracy.
@@ II Related Works
II-A Simulatenous Localization and Mapping
Metric-Semantic SLAM. The most referenced LiDAR SLAM pipelines include LOAM [5], LIO-SAM [6], BALM [7], Fast-LIO [8], [9]. All the above techniques utilize low-level geometric features in the environment to estimate the pose and map of the robot, due to which such techniques can be limited in their accuracy when exploring large and complex indoor environments. Additionally several other SLAM techniques exist in the literature that utilize high-level semantic features additional to geometric features to improve the environment understanding and the map and pose accuracy. Some of such techniques include LeGO-LOAM [10], SA-LOAM [11], SegMap [12], SUMA++ [13]. LeGO-LOAM [10] utilizes different planar semantics in the environment like ground plane to improve the map and pose estimate. While SA-LOAM [11] utilizes high-level semantics to improve the loop closure accuracy of the underlying metric SLAM, SegMap [12] utilizes learned high-level descriptors from the environment to perform robust localization with respect to the high-level descriptors. SUMA++ [13] completely segments the environment in different semantic features to perform object level semantic SLAM removing dynamic entities from the scene.
@@ II-A Simulatenous Localization and Mapping
Hierarchical SLAM. To overcome the inherent problem of improving the computation with increasing map size, works like [14], [15] exploit the methods to compress the graph into different sub-graphs to maintain real-time performance without the loss of map and pose accuracy. [14] present a technique grouping nodes into different sub-graphs based on a simple distance based criteria. While [15] present an information-theoretic approach for factor graph compression where laser scans measurements and their corresponding robot poses are removed such that the expected loss of information with respect to the current map is minimized. In [16] authors present a hierarchical continuous time SLAM algorithm dividing it into local sub-graphs aggregating measurements from 3D LiDAR, the generated sub-graphs are aligned with each other using local surfel based registration techniques. In the above technique authors also use distance based heuristic to create different sub-graphs for the optimization problem. [17] propose a local and global hierarchical optimization technique using sub-map strategies similar to [18]. They generate local sub-maps at given distance intervals performing local optimization and then connecting different local sub-maps using global registration to perform efficient global optimization. The authors in [17] also choose a heuristic based stragety to generate the local and global maps.
@@
Recently, 3D scene graphs have shown great potential in representing the environment in a more meaningful and compact manner [19, 20, 21, 22, 23]. Additionally methods such as, [2], [3], [4] tightly couple the 3D scene graph with SLAM graphs exploiting the hierarchy in the environment to improve the final pose and map accuracy. [2] focuses on generation of 3D scene graph mainly using RBG-D cameras incorporating different hierarchies of objects, places, rooms and buildings, furthermore it utilizes the hierarchy to perform a top-down and bottom-up search improving the search of loop closure candidates, thus improving the final map accuracy. [3] presents a 3D scene graph for outdoor environment dividing it in hierarchies such as lanes, landmarks, intersections and environment. They factor different connections of the hierarchical graph as a factor graph, continuously optimizing the pose and map. [24, 4] represent an indoor scene as a four-layered optimizable factor graph dividing it into layers of keyframes, walls, rooms and floors, simultaneously optimizing all the layers to obtain improved map and pose accuracy.
@@
Thus in this work, we apply the concept of hierarchical SLAM to 3D scene graphs. The 3D scene graph structure allows to generate meaningful hierarchical SLAM graph instead of using distance-time based heuristics. This hierarchical SLAM graph improves the management of generated map elements to scale for large scale environments, while maintaining the pose and map accuracy with real-time performance.
@@ III-A Background
Situational Graphs. We build over S-Graphs+ [4], that estimates a layered graph representation of the environment from 3D LiDAR readings. The four layers of the graph, illustrated in Fig. 2, are as follows. The Keyframes Layer, containing robot poses recorded at different distance-time intervals. The Walls Layer, representing the walls in the environment as planar surfaces connected to the keyframes from which they were observed. The Rooms Layer, modeling two-wall and four-wall rooms in the environment, connecting in this manner the previous walls layer. Finally, the Floors Layer that represents the different floors in the building, connecting them with the underlying rooms layer.
@@ III-B System Architecture
Fig. 2 shows the complete architecture of our approach. The front-end implement plane and room segmentation methods, as in [4], with the addition of a floor segmentation module able to detect changes in the floor levels by measuring changes in the slope of the trajectory of keyframes. In practice, floor segmentation will be highly relevant to perform loop closure only within the same floor. This is a reasonable strategy, as local sensor readings are commonly limited to a single floor and connecting staircases. Additionally, it removes a high rate of false positives, caused by aliasing between floor levels.
@@ IV-A Floor Segmentation
Floor Center. The floor segmentation module extracts the center of a current floor level from the largest wall planes currently extracted in a scene. Each time the pipeline is run, it creates a default floor node with the center placed at the origin of frame . It then utilizes the information from all walls at the current floor level to create a sub-category of wall planes where . The walls sub-categories consist of vertical planar surfaces and vertical planar surfaces based on their normal orientation. After receiving a plane set, it computes the widths between all -direction planes and similarly for -direction planes as in [4].
@@ IV-C Room Keyframe Segmentation
In order to perform room-level local optimization (explained in Section. V) we need to identify keyframes which are bounded within a four-wall room. We exploit the hierarchy within a room and its corresponding four-walls to identify the keyframes set belonging to a room . To compute if a keyframe lies within a room, we first compute the vector , where is the translation component of the keyframe and is the point on one of the planes belonging to the room. We then evaluate the dot product of , where is the normal orientation of the plane. Keyframe positions with positive are considered to be bounded within the walls of the room.
@@ V Back-End
The global state of the robot defined in III-B is optimized in three different stages exploiting the hierarchy of the graph. We mainly detail the optimization strategies, where relies the novelty of the paper. Details on the formulation of the cost function can be found in [4].
@@ V-A Local Optimization
As a robot navigates through a scene we generate specific keyframe nodes and its observations thus incorporating new nodes and its edges within the graph. Every time the graph increases in size we perform a local optimization over a subset of a graph called local graph choosing a window of last keyframes. In order to incorporate other hierarchical connected layers from the graph in the local optimization, we first check the connections of with the walls layer, incorporating all the wall nodes and their edges with keyframes . Given all the wall within the window, if the entire set or a subset of it have connections to two-wall or four-wall rooms we incorporate these rooms and within the local optimization. Furthermore, we incorporate into the local graph the corresponding the floor node the robot is currently navigating. In order to maintain consistency of the graph during optimization, we follow the strategy proposed in [26] to fix keyframes that are outside the current local optimization window but observe any of the walls inside the local window. The state is be optimized as:
where are the cost functions for the unfixed keyframes, walls and rooms, is the cost of the fixed keyframes and is the cost of the floor node currently being explored. Fig. 4 shows the local optimization strategy, with orange color keyframes along with the connected layers being optimized while the fixed keyframes highlighted in blue.
@@ V-B Floor-level Global Optimization
Floor-level global optimization is performed in mainly two situations, specifically, when a loop closure is found between the keyframes of a floor-level, and when a duplicate wall is found when a room is re-detected as explained in [4]. Our main motivation for this strategy is as follows. When a new floor-level is explored, and a loop closure is found between two keyframes of such floor-level, this loop closure does not affect the keyframes or other connected layers of the preceding floor-levels. The same reasoning applies when duplicate walls are found for a detected room.
When executing floor-level global optimization for floor-level , we create a subset of the graph containing floor node , keyframe nodes and its edges with the neighboring keyframes, then incorporate all wall nodes with their edges to keyframe nodes . Floor-level global optimization incorporates all previously identified loop closure (relative pose) edges between the keyframes of that floor-level. We also incorporate four-wall and two-wall room nodes with their edges to the underlying walls, finally also incorporating the edges between these rooms and the current floor node. To maintain consistency of the graph during optimization, we fix a keyframe which can either be the initial keyframe when the robot starts navigating, or it can be the last keyframe from the previous visited floor-level, this assures that the current floor-level map estimates do not diverge away from the previous floor-level map (See Fig. 5). The state using floor-level global optimization is optimized as:
@@
As a relevant effect of floor-level global optimization, if a robot revisits a given floor-level and finds a suitable loop closure candidate, the floor-level global optimization incorporates all the nodes and edges not only from the current floor-level but also from all the previous visited ones. This ensures the correction of all the accumulated errors from the revisited floor-level, as well as the previously visited ones.
@@ V-C Room-level Local Optimization
where:
is the information matrix of the original edge .
is the resulting information matrix for the new edge .
The new edge is added back to the global graph edges:
@@ VI-A Methodology
We validate our algorithm on different real indoor environments from university buildings to construction sites, validating it with different state-of-the-art SLAM algorithms including the baseline S-Graphs+ [4]. Our datasets are recorded using either using Velodyne VLP-16 or Ouster OS-1 64 3D LiDAR, with the odometry generated either using robot encoders or FAST-LIO2 [9]. We perform the experiments in real in-house datasets dividing them into single-floor or multi-floor datasets.
@@
Single-Floor Dataset. The single-floor dataset consists of small-scale indoor environments with the robot navigating around one single floor. The main motivation for these experiments is to fairly validate the accuracy and cost reduction of our S-Graphs 2.0 against baselines, as most baselines accumulate significant drift or incur in false loop closures for multi-floor scenes. The first two experiments, C1F1 and C1F2, are performed on two different floors of a construction site. Additionally, C2F0, C2F1, and C2F2 consist of three different floors of an ongoing construction site combining four individual houses. C3F1, and C3F2 are two different floors two combined houses. To validate the accuracy of the given approaches we use the RMSE of the generated 3D map against the ground truth 3D map. Furthermore, we also report the computational time of our approach against the most accurate baseline S-Graphs+ [4].
@@
Multi-Floor Dataset. The multi-floor dataset is aimed at validating the accuracy of our proposed approach with baselines in presence larger trajectories covering multiple floors. It consists of five datasets ranging from BE1 and BE2 which are university buildings where the robot navigates the three and two floors respectively. LC1 consists of a university building with long corridors and robot traversing two floors with both the floors resembling each other. CS1 and CS2are buildings in construction with robot traversing all the three floors. Due to absence of ground truth, to validate the accuracy of each approach we use the Mean Map Entropy (MME) presented in [27] to measure the sharpness of the generated map. Additionally, to evaluate the accuracy of the algorithms to properly segment multiple floors, we perform an evaluation of Intersection over Union (IoU) of histogram of occurrences of values per floor-level from the generated map versus the percentage ground truth time spent per floor-level.

### v3

@@ Abstract
The hierarchical nature of 3D scene graphs aligns well with the structure of man-made environments, making them highly suitable for representation purposes. Beyond this, however, their embedded semantics and geometry could also be leveraged to improve the efficiency of map and pose optimization, an opportunity that has been largely overlooked by existing methods. We introduce Situational Graphs 2.0 (S-Graphs 2.0), that effectively uses the hierarchical structure of indoor scenes for efficient data management and optimization. Our approach builds a four-layer situational graph comprising Keyframes, Walls, Rooms, and Floors. Our first contribution lies in the front-end, which includes a floor detection module capable of identifying stairways and assigning floor-level semantic relations to the underlying layers (Keyframes, Walls, and Rooms). Floor-level semantics allows us to propose a floor-based loop closure strategy, that effectively rejects false positive closures that typically appear due to aliasing between different floors of a building. Our second novelty lies in leveraging our representation hierarchy in the optimization. Our proposal consists of: (1) local optimization over a window of recent keyframes and their connected components across the four representation layers, (2) floor-level global optimization, which focuses only on keyframes and their connections within the current floor during loop closures, and (3) room-level local optimization, marginalizing redundant keyframes that share observations within the room, which reduces the computational footprint. We validate our algorithm extensively in different real multi-floor environments. Our approach shows state-of-the-art accuracy metrics in large-scale multi-floor environments, estimating hierarchical representations up to faster, in average, than competing baselines. Our code is open-sourced at: https://github.com/snt-arg/lidar_situational_graphs
I Introduction
@@
Simultaneous Localization and Mapping (SLAM), that has typically focused on geometric aspects, should also enable a robot with a high-level understanding of its environment [1]. The combination of 3D scene graphs with traditional SLAM graphs have recently sparked great interest given its compact and intuitive approach to representing an environment as a hierarchy of layers [2], [3]. In this direction, Situational Graphs (S-Graphs) [4] model the environment as a four-layered optimizable graph composed of of Keyframes, Walls, Rooms and Floors. As the robot explores large environments, however, the size of the graph grows, along with the its computational footprint. Essentially, while S-Graphs and 3D scene graphs are motivated by the need of higher-level representations, none of them have yet tailored their optimization to the specifics of these hierarchical representations. Leveraging them will result in a lower computational footprint, enabling real-time graph optimization for larger maps, which will result finally in enhanced accuracy.
@@ II Related Works
II-A Simultaneous Localization and Mapping
Metric-Semantic SLAM. The most referenced LiDAR SLAM pipelines include LOAM [5], LIO-SAM [6], BALM [7], Fast-LIO [8], [9]. All the above techniques utilize low-level geometric features in the environment to estimate the pose and map of the robot, due to which such techniques can be limited in their accuracy when exploring large and complex indoor environments. Additionally several other SLAM techniques exist in the literature that utilize high-level semantic features additional to geometric features to improve the environment understanding and the map and pose accuracy. Some of such techniques include LeGO-LOAM [10], SA-LOAM [11], SegMap [12], SUMA++ [13]. LeGO-LOAM [10] utilizes different planar semantics in the environment like ground plane to improve the map and pose estimate. While SA-LOAM [11] utilizes high-level semantics to improve the loop closure accuracy of the underlying metric SLAM, SegMap [12] utilizes learned high-level descriptors from the environment to perform robust localization with respect to the high-level descriptors. SUMA++ [13] completely segments the environment in different semantic features to perform object level semantic SLAM, removing dynamic entities from the scene.
@@ II-A Simultaneous Localization and Mapping
Hierarchical SLAM. To overcome the inherent problem of improving the computation with increasing map size, works like [14], [15] exploit the methods to compress the graph into different sub-graphs to maintain real-time performance without the loss of map and pose accuracy. [14] present a technique grouping nodes into different sub-graphs based on a simple distance based criteria. While [15] present an information-theoretic approach for factor graph compression where laser scans measurements and their corresponding robot poses are removed such that the expected loss of information with respect to the current map is minimized. In [16] authors present a hierarchical continuous time SLAM algorithm dividing it into local sub-graphs aggregating measurements from 3D LiDAR, the generated sub-graphs are aligned with each other using local surfel-based registration techniques. In the above technique authors also use distance-based heuristic to create different sub-graphs for the optimization problem. [17] propose a local and global hierarchical optimization technique using sub-map strategies similar to [18]. They generate local sub-maps at given distance intervals performing local optimization and then connecting different local sub-maps using global registration to perform efficient global optimization. The authors in [17] also choose a heuristic based stragety to generate the local and global maps.
@@
Recently, 3D scene graphs have shown great potential in representing the environment in a more meaningful and compact manner [19, 20, 21, 22, 23]. Additionally methods such as, [2], [3], [4] tightly couple the 3D scene graph with SLAM graphs exploiting the hierarchy in the environment to improve the final pose and map accuracy. [2] focuses on the generation of a 3D scene graph mainly using RGB-D cameras, incorporating different hierarchies of objects, places, rooms and buildings, furthermore it utilizes the hierarchy to perform a top-down and bottom-up search improving the search of loop closure candidates, thus improving the final map accuracy. [3] presents a 3D scene graph for the outdoor environment, dividing it into hierarchies such as lanes, landmarks, intersections, and environment. They factor different connections of the hierarchical graph as a factor graph, continuously optimizing the pose and map. [24, 4] represent an indoor scene as a four-layered optimizable factor graph dividing it into layers of keyframes, walls, rooms, and floors, simultaneously optimizing all the layers to obtain improved map and pose accuracy.
@@
Thus in this work, we apply the concept of hierarchical SLAM to 3D scene graphs. The 3D scene graph structure allows to generate a meaningful hierarchical SLAM graph instead of using distance-time based heuristics. This hierarchical SLAM graph improves the management of generated map elements to scale for large scale environments, while maintaining the pose and map accuracy with real-time performance.
@@ III-A Background
Situational Graphs. We build on S-Graphs+ [4], that estimates a layered graph representation of the environment from 3D LiDAR. The four layers of the graph, illustrated in Fig. 2, are as follows. The Keyframes Layer contains robot poses at different distance-time intervals. The Walls Layer represents the walls in the environment as planar surfaces connected to the keyframes from which they were observed. The Rooms Layer models two-wall and four-wall rooms, connecting in this manner the previous walls layer. Finally, the Floors Layer represents the different floors in the building, connecting them with the underlying rooms layer.
@@ III-B System Architecture
Fig. 2 shows the complete architecture of our approach.The front-end implements plane segmentation as in [4] and room segmentation as in [25], with the addition of a floor segmentation module able to detect changes in the floor levels by measuring changes in the slope of the trajectory of keyframes. In practice, floor segmentation will be highly relevant to perform loop closure only within the same floor. This is a reasonable strategy, as local sensor readings are commonly limited to a single floor and connecting staircases. Additionally, it removes a high rate of false positives, caused by aliasing between floor levels.
@@ IV-A Floor Segmentation
Floor Center. The floor segmentation module extracts the center of a current floor level from the largest wall planes currently extracted in a scene. Each time the pipeline is run, it creates a default floor node with the center placed at the origin of frame . It then utilizes the information from all walls at the current floor level to create a sub-category of wall planes ( represents a set of planes) where . The walls sub-categories consist of vertical planar surfaces and vertical planar surfaces based on their normal orientation. After receiving a plane set, it computes the widths between all -direction planes and similarly for -direction planes as in [4].
@@ IV-C Room Keyframe Segmentation
In order to perform room-level local optimization (explained in Section V) we need to identify keyframes which are bounded within a four-wall room. We exploit the hierarchy within a room and its corresponding four-walls to identify the keyframes set belonging to a room . To compute if a keyframe lies within a room, we first compute the vector , where is the translation component of the keyframe and is a randomly selected point from the -th plane of the room, where corresponds to each of the four planes belonging to the room. We then evaluate the dot product of , where is the normal orientation of the plane. Keyframe positions with positive are considered to be bounded within the walls of the room.
@@ V Back-End
The global state of the robot defined in Section III-B is optimized in three different stages exploiting the hierarchy of the graph. We mainly detail the optimization strategies, where relies the novelty of the paper. Details on the formulation of the cost function can be found in [4].
@@ V-A Local Optimization
As a robot navigates through a scene we generate specific keyframe nodes and its observations thus incorporating new nodes and its edges within the graph. Each time the graph expands, we perform local optimization on a subset of the complete graph called the local graph, which consists of a sliding window containing the most recent keyframes. In order to incorporate nodes and edges belonging to other hierarchical layers of the graph into local graph, we first check the connections between the keyframes and the wall layer, incorporating all connected wall nodes and their corresponding edges. For all wall nodes within the optimization window, we check whether the entire set or a subset belongs to two-wall rooms or four-wall rooms . If such connections exist, we incorporate these room nodes into the local optimization.
Furthermore, we incorporate the corresponding floor node that the robot is currently navigating into the local graph. To maintain graph consistency during optimization, we adopt the strategy proposed in [27], fixing keyframes that lie outside the current optimization window but observe walls within the local window. The state to be optimized is defined as:
where are the cost functions for the unfixed keyframes, walls and rooms respectively , is the cost of the fixed keyframes and is the cost of the floor node currently being explored. Fig. 4 shows the local optimization strategy, with orange color keyframes along with the connected layers being optimized while the fixed keyframes are highlighted in blue.
@@ V-B Floor-level Global Optimization
Floor-level global optimization is performed in mainly two situations, specifically, when a loop closure is found between the keyframes of a floor-level, and when a duplicate wall is found when a room is re-detected as explained in [4]. Our main motivation for this strategy is as follows. When a new floor-level is explored, and a loop closure is found between two keyframes of such floor-level, this loop closure does not affect the keyframes or other connected layers of the preceding floor-levels. The same reasoning applies when duplicate walls are found for a detected room. When executing floor-level global optimization for floor-level , we create a subset of the graph containing floor node , keyframe nodes and its edges with the neighboring keyframes, then incorporate all wall nodes with their edges to keyframe nodes . Floor-level global optimization incorporates all previously identified loop closure (relative pose) edges between the keyframes of that floor-level. We also incorporate four-wall and two-wall room nodes with their edges to the underlying walls, finally also incorporating the edges between these rooms and the current floor node. To maintain consistency of the graph during optimization, we fix a keyframe which can either be the initial keyframe when the robot starts navigating, or it can be the last keyframe from the previous visited floor-level, this assures that the current floor-level map estimates do not diverge away from the previous floor-level map (See Fig. 5). The state using floor-level global optimization is optimized as:
@@
As a relevant effect of floor-level global optimization, if a robot revisits a given floor-level and finds a suitable loop closure candidate, the floor-level global optimization incorporates all the nodes and edges not only from the current floor-level but also from all the previous visited ones. This approach also enables loop closures spanning multiple floors through staircases, ensuring comprehensive error correction across the entire multi-floor structure. This ensures the correction of all the accumulated errors from the revisited floor-level, as well as the previously visited ones.
@@ V-C Room-level Local Optimization
where is the information matrix of the original edge and is the resulting information matrix for the new edge . The new edge is added back to the global graph edges:
@@ VI-A Methodology
We validate our algorithm on different real indoor environments (single-floor and multi-floor) from university buildings to construction sites, validating it with different state-of-the-art SLAM algorithms, including the baseline S-Graphs+ [4]. Our datasets are recorded using Velodyne VLP-16 or Ouster OS-1 64 3D LiDAR, with odometry from robot encoders or FAST-LIO2 [9]. These sensors demonstrate that our algorithm is sensor-agnostic. Despite differing beam counts, both sensors are downsampled to uniform resolution for processing. Dynamic objects such as humans are present in many of these datasets; while handling dynamic objects falls outside the scope of this paper, the reported results reflect their presence as we neither explicitly include nor exclude them from our analysis.
@@
Single-Floor Dataset. The single-floor dataset consists of small-scale indoor environments with the robot navigating around one single floor. The main motivation for these experiments is to fairly validate the accuracy and cost reduction of our S-Graphs 2.0 against baselines, as most baselines accumulate significant drift or incur in false loop closures for multi-floor scenes. The first two experiments, C1F1 and C1F2, are performed on two different floors of a construction site. Additionally, C2F0, C2F1, and C2F2 consist of three different floors of an ongoing construction site combining four individual houses. C3F1, and C3F2 are two different floors two combined houses. For validation, due to the absence of ground truth trajectory metrics, we use ground truth point clouds, extracted from available Building Information Models (BIM), to report pointcloud RMSE. Furthermore, we also report the computational time of our approach against the most accurate baseline S-Graphs+ [4].
@@
Multi-Floor Dataset. The multi-floor dataset is aimed at validating the accuracy of our algorithm in the presence of larger trajectories covering multiple floors. It consists of five datasets, BE1 and BE2 are university buildings where the robot navigates the three and two floors, respectively. LC1 is a university building with long corridors and a robot traversing two floors. CS1 and CS2 are buildings under construction with a robot traversing all three floors. Due to the absence of ground truth, to validate the accuracy of each approach, we use the Mean Map Entropy (MME) presented in [28] to measure the sharpness of the generated map. Additionally, to assess algorithmic accuracy in multi-floor segmentation, we evaluate the Intersection over Union (IoU) between the histogram of z-value occurrences per floor level in the generated map and the percentage of ground truth time spent per floor level. In this evaluation, we recorded the time the robot spent on each floor level, which was subsequently converted into percentage data for that floor. We also documented the maximum and minimum heights for each floor level, establishing the ground truth bins for the datasets. We then utilized the z-values from the 3D maps generated by our algorithms and baselines to compute the IoU of these bins against the ground truth bins. While IoU is an indirect metric, it captures if the SLAM algorithm correctly tracks the robot’s position. The vertical distribution of mapped points should match the time spent on each floor. When the algorithm incorrectly identifies floors, points are mapped at incorrect heights, resulting in clear mismatches in the z-distribution. Thus, high IoU indicates good floor-level tracking throughout the mapping process, providing meaningful validation despite the lack of ground truth poses.
VI-B Results and Discussions
Single-Floor Dataset. The results of the single-floor experiments, in Tab. I, show that our S-Graphs 2.0 is able to provide final map accuracy similar to the second best baseline S-Graphs+ with an average accuracy improvement of . However, the benefits of the hierarchical optimization of our S-Graphs 2.0 clearly stand out in Tab. V, where it has an average cost reduction of to the closest baseline. The benefits of the hierarchical optimization become even more clear for longer sequences. See, for example, how in C2F2, C3F2 there is slight decrease in the accuracy with respect to S-Graphs+ ( and respectively),but a large reduction the computation time ( and respectively). Summing up, in these single-floor datasets, we show that the optimization strategy in S-Graphs 2.0 maintains the estimation accuracy while significantly reducing the computational cost. We also compared our algorithm and other baselines on the publicly available Hilti SLAM datasets [29]. For fair comparison, we use the same odometry source fast lio [9] for all. Tab. II reports the ATE for the easy (E04, E05) and medium (E06, E14) difficulty categories. We used Our S-Graphs 2.0 consistently outperforms the baseline in all sequences. Note that none of the methods converged on the hard category sequences due to their extreme difficulty.